two facing pages product of page # 210 what are page numbers?


Answer by unlockmath(1688)

Hello,
Let x be one page and (x+1) be the next so we can set up the following equation:
x(x+1)=210
Expand this:
x^2+x=210
Subtract 210 from both sides to get:
x^2+x-210=0
This can be factored as:
(x+15)(x-14)=0
Solve for x:
x=-15
x=14
14 is the sensible answer. So the pages are:
14 & 15
Make sense?
RJ
www.math-unlock.com